OpenStudy (huangtina):

What is the fastest method to calculate a non-perfect square root without using a calculator?

OpenStudy (anonymous):

Probably something like newton's method.

OpenStudy (anonymous):

Suppose you are trying to find the square root of \(c\). This is basically the same as finding the roots of \(f(x) = x^2-c\).

OpenStudy (solomonzelman):

Wouldn't you still then have to use a calculator?

OpenStudy (anonymous):

The derivative of this is: \( f'(x) = 2x \). Newton's method says start with some guess \(x_0\), the following number will be closer to the square root: \[ x_1 = x_0-\frac{f(x_0)}{f'(x_0)} = x_0 - \frac{x_0^2-c}{2x_0} \]

OpenStudy (anonymous):

You can keep going until you think you're close enough.

OpenStudy (solomonzelman):

A second degree taylor polynomial would be fairly close usually.

OpenStudy (solomonzelman):

But, this is an algebra question. (not calculus, so offering differentiation to solve it??)

OpenStudy (solomonzelman):

You can estimate c between two perfect squares. (I have seen that technique when I was helping my relative with math lessons on connections academy)

OpenStudy (huangtina):

This is for my student who is 11 years old....so maybe something easier? Or is estimate and calculate and estimate the only real way?

OpenStudy (anonymous):

For example, suppose I want the square root of \(13\). I guess \(4\).\[ x_1 = 4 - \frac{4^2-13}{2(4)} = 4 - \frac{3}{8} = \frac{29}{8} = 3.625 \]Then I can do this again: \[ x_2 = 3.625 - \frac{3.625^2-13}{2(3.625)} \]

OpenStudy (anonymous):

for finding perfect square of a no. which wasn't be a perfect square, you need to know minimum of square root of first 20 natural no. let you need to find sqrt 80.\[\sqrt{80}\]than first calculate LCM of this,now simply wright\[\sqrt{80}=\sqrt{2*2*2*2*5}\]now wright this as \[2*2\sqrt{5}\]now simply put the value of \[\sqrt{5}\] and you get required answer.

OpenStudy (huangtina):

Hmmm, wio that is not too bad.

OpenStudy (anonymous):

Next iteration I get: \[ x_2 = \frac{1673}{464} \approx 3.6056 \]Note that \[ \sqrt{13} \approx 3.60555 \]

OpenStudy (anonymous):

Another method, that could be a bit easier, is to memorize the square root of prime numbers up to 25 or so, then to factor as newtonson suggests.

OpenStudy (anonymous):

This is the method that engineers used before calculators.

OpenStudy (huangtina):

You still really need a calculator. I will review newtonson

OpenStudy (anonymous):

Except, they had a reference table for square roots of primes.

OpenStudy (anonymous):

Typically math classes that don't allow calculators will allow you to keep the square root of a prime, and if not, then they expect you to have memorized the square root of the prime.

OpenStudy (huangtina):

Thanks everybody.